### v1

# The Impact of Auxiliary Patient Data on Automated Chest X-Ray Report Generation and How to Incorporate It

Abstract
This study investigates the integration of diverse patient data sources into multimodal language models for automated chest X-ray (CXR) report generation. Traditionally, CXR report generation relies solely on CXR images and limited radiology data, overlooking valuable information from patient health records, particularly from emergency departments. Utilising the MIMIC-CXR and MIMIC-IV-ED datasets, we incorporate detailed patient information such as aperiodic vital signs, medications, and clinical history to enhance diagnostic accuracy. We introduce a novel approach to transform these heterogeneous data sources into embeddings that prompt a multimodal language model, significantly enhancing the diagnostic accuracy of generated radiology reports. Our comprehensive evaluation demonstrates the benefits of using a broader set of patient data, underscoring the potential for enhanced diagnostic capabilities and better patient outcomes through the integration of multimodal data in CXR report generation.
The Impact of Auxiliary Patient Data on Automated Chest X-Ray Report Generation and How to Incorporate It
Aaron Nicolson, Shengyao Zhuang, Jason Dowling, & Bevan Koopman Australian e-Health Research Centre, CSIRO Health and Biosecurity, Brisbane, Australia aaron.nicolson@csiro.au
1 Introduction
Figure 1: The patient data from MIMIC-IV-ED associated with a CXR exam from MIMIC-CXR. This includes the exam’s images, the corresponding radiology report, and the associated image metadata. The findings and impression sections of the radiology report form the ground truth for CXR report generation. ED-specific data, such as medicine reconciliation and aperiodic vital signs, is also available for the patient.
Chest X-ray (CXR) exams, which consist of multiple images captured during an imaging session, are essential for diagnosing and managing a wide range of conditions, playing a significant role in patient care. Radiologists interpret these exams and produce a written report with their findings. However, prompt reporting is hindered by a multitude of issues, including high patient volumes and limited availability of radiologists (Bailey et al., 2022).
Machine learning for automated CXR report generation is a promising solution that has garnered significant attention in the literature (Jones et al., 2021). By leveraging multimodal language models, exams can be rapidly interpreted and reported, potentially providing quick and reliable diagnostic insights crucial for decision-making, such as triaging patients. Models are often trained to generate the findings and impression sections of a radiology report (Figure 1), where the former details the interpretation of a patient’s exam and the latter summarises the most important findings. Potential benefits include enhanced radiologist effectiveness, a reduced workload, alleviation of the burden of report writing, and improved patient outcomes (Shen, 2021; Irmici et al., 2023).
Early methods for CXR report generation produced a separate report for each image within an exam Wang et al. (2018). Later methods improved on this by considering all images of an exam to generate a single report Miura et al. (2021); Nicolson et al. (2024a), and incorporating prior exams for a patient Wu et al. (2022); Nicolson et al. (2024a). Additionally, including the reason for conducting the exam (the indication section in Figure 1) offered a further improvement Nguyen et al. (2023). This indicates that CXR report generation could benefit from the inclusion of a more comprehensive set of patient data.
Modern patient record systems are another rich source of patient data, containing detailed information that may be valuable for CXR report generation. However, (1) the utility of this data has not been empirically investigated, and (2) it is unclear how to harmonise this heterogeneous data into a unified multimodal language model. This paper aims to address these two points. To achieve this, we combine CXR exams from MIMIC-CXR (Johnson et al., 2019) with emergency department (ED) patient records from MIMIC-IV-ED (Johnson et al., 2023). This means that for a single exam, a wide variety of multimodal data is available, as shown in Figure 1. From MIMIC-CXR, we utilise the images, their metadata, and several sections of the radiology report. Notably, incorporating the comparison or history section is a novel approach in the literature. From MIMIC-IV-ED, we investigate triage information, aperiodic vital signs, medications, and other data to provide a wider clinical context.
We explore combining these sources of patient data as patient embeddings to prompt a multimodal language model. We demonstrate that complementary information from different data sources can improve the diagnostic accuracy of CXR report generation. To achieve this, we develop methods to transform tabular and aperiodic time series data into embeddings that can be used alongside token and image embeddings. We evaluate our model on MIMIC-CXR exams with accompanying patient data from MIMIC-IV-ED, using metrics shown to closely correlate with radiologists’ assessments of reporting (Yu et al., 2023). The main contributions of this work are:
  • An investigation into how patient data impacts CXR report generation, focusing on the effects of specific data sources, such as medications and vital signs.
  • An empirical evaluation demonstrating that using multiple patient data sources — from a patient’s CXR exams and their ED record — significantly improves diagnostic accuracy.
  • Introducing methods to convert multimodal patient data into embeddings for a language model, including numerical, categorical, free text, temporal, and image data.
  • A release of dataset splits based on MIMIC-CXR and MIMIC-IV-ED, linking patient exams with their associated ED records (available as a Hugging Face dataset). This, along with our code repository and Hugging Face checkpoint can be found at: https://anonymous.4open.science/r/anon-D83E, enabling others to experiment with new methods for multimodal patient data.
2 Background and Related Work
There is evidence to suggest that incorporating more patient data improves diagnostic accuracy in radiology reporting. Initial improvements came from using multiple images per exam, like EMNLI, which often includes complementary frontal and lateral views of the patient Miura et al. (2021); Gaber et al. (2005). Methods such as CXRMate enhance diagnostic accuracy by incorporating a patient’s prior exams to identify changes over time Nicolson et al. (2024a); Wu et al. (2022); Kelly (2012). Including the indication section of the radiology report to provide clinical context also provides an improvement Nguyen et al. (2023). This trend indicates that providing more comprehensive patient data improves diagnostic accuracy, which we investigate in this work.
ED records contain a myriad of data, including vital signs such as respiratory rate, temperature, and blood pressure, which can aid in the identification of various diseases. A high respiratory rate and low blood oxygen saturation are indicative of conditions that compromise pulmonary function, such as pulmonary embolism. Similarly, an elevated body temperature is suggestive of an infectious process, such as pneumonia or tuberculosis. Incorporating such data into a CXR report generator could help corroborate subtle radiographic signs typical of these infections. Our findings demonstrate that patient data from the ED can indeed enhance CXR report generation.
Recent advancements in integrating multimodal patient data have enhanced diagnostic and predictive healthcare capabilities. A study showed that a Transformer encoder combining imaging and non-imaging data outperformed single-modality models, diagnosing up to 25 conditions with higher AUC scores Khader et al. (2023b). Similarly, the MeTra architecture, which integrates CXRs and clinical parameters, demonstrated superior performance in predicting ICU patient survival compared to using either CXRs or clinical data alone Khader et al. (2023a). ETHOS, using a zero-shot learning approach, outperformed single-modality models in predicting inpatient mortality, ICU length of stay, and readmission rates Renc et al. (2024). These studies highlight the importance of multimodal data for improved healthcare analytics. Our work demonstrates that incorporating a comprehensive set of multimodal patient data enhances CXR report generation.
Recent advancements in multi-task learning have significantly improved biomedical models by leveraging shared knowledge. Med-PaLM M, a generalist biomedical model, excels in multiple tasks including classification, question answering, visual question answering (VQA), report summarisation, report generation, and genomic variant calling, using diverse input modalities like images, text, and genomics. It often outperforms specialised models, demonstrating superior performance and generalisation (Tu et al., 2024).
Similarly, MIMIC-CXR has been leveraged for multi-task learning with models like MedXChat, which integrates instruction-tuning and Stable Diffusion to perform CXR report generation, VQA, and report-to-CXR generation, outperforming other LLM multi-task learners (Yang et al., 2023). RaDialog, another LLM-based method, combines visual features and pathology findings to generate accurate radiology reports and support interactive tasks, significantly improving clinical efficacy. CXR-LLaVA, a multimodal LLM integrating a vision transformer with a language model, outperformed models like GPT-4 Vision and Gemini Pro Vision in CXR report generation Lee et al. (2024).
Determining the state-of-the-art CXR report generation model can be challenging due to the unavailability of some models and the lack of comparison to recent methods. The 2024 Shared Task on Large-Scale Radiology Report Generation (RRG24) aimed to address this by benchmarking models on a common leaderboard. The winning model, CXRMate-RRG24 Nicolson et al. (2024b), a derivative of CXRMate, emerged as a strong contender for state-of-the-art. In this work, we compare our model to established models (e.g., EMNLI) and recent benchmarks (e.g., CXRMate-RRG24, CXRMate, CXR-LLaVA, MedXChat, and RaDialog). We ensure a fair comparison by using available code or obtaining generated reports directly from the authors. Our findings indicate our model produces significantly better results than these models.
Figure 2: Multimodal language model for CXR report generation. The patient data embeddings prompt the decoder to generate the findings and impression sections of a radiology report.
3 Dataset
We construct a dataset of 46 1064610646\,10646 106 patients by linking individual patient information from two separate sources: (1) CXR exams from MIMIC-CXR and (2) emergency records from MIMIC-IV-ED. Thus we consider MIMIC-CXR exams that occurred during an ED stay from MIMIC-IV-ED. Both datasets are publicly available and originate from the Beth Israel Deaconess Medical Center in Boston, MA.
MIMIC-CXR was formed by first extracting patient identifiers for exams performed in the ED between 2011–2016, and then extracting all exams for this set of patients from all departments between 2011–2016. Each exam includes a semi-structured free-text radiology report (Figure 1) that describes the radiological findings of the images, written by a practising radiologist contemporaneously during routine clinical care. All images and reports were de-identified to protect privacy. Sections from the radiologist reports were extracted using a modification [^1] https://anonymous.4open.science/r/anon-D83E of the official text extraction tool [^2] https://github.com/MIT-LCP/mimic-cxr/tree/master/txt in order to obtain the findings, impression, indication, history, and comparison sections.
MIMIC-IV-ED consists of de-identified data from ED stays between 2011–2019. The data was converted into a denormalised relational database with six primary tables: ED stays, diagnosis, medicine reconciliation, medicine administration, triage, and aperiodic vital signs. We do not consider the diagnosis table in this work, as it indicates the outcome of a patient’s ED stay. The patients of MIMIC-CXR can be linked to MIMIC-IV-ED via an identifier, allowing an ED specific dataset to be formed.
Example tables for a patient’s exam are shown in Figure 1. The dataset was formed by extracting a patient’s exams whose times (formed by the ‘StudyDate’ and ‘StudyTime’ columns of the metadata table) occurred within the ‘intime’ and ‘outtime’ of one of their ED stays. [^3] Exam 59128861 was removed as it overlapped with two separate ED stays for the patient. Exams with either a missing findings or impression section were not considered. Using the official splits of MIMIC-CXR, this gave a train/validation/test split of 45 5274552745\,52745 527/343/236 patients, 76 3987639876\,39876 398/556/958 exams, and 151 818151818151\,818151 818/1 13711371\,1371 137/1 81218121\,8121 812 CXRs. Each of these exams had one ED stay and triage row; 53% had at least one medicine reconciliation row with up to 106 rows; 62% had at least one vital signs row with up to 69 rows; and 37% had at least one medication administration row with up to 52 rows. Exams had an indication section 66% of the time with a maximum of 75 words, a history section 34% of the time with a maximum of 74 words, and a comparison section 97% of the time with a maximum of 129 words. Only one exam had both an indication and a history section.
Figure 3: Proposed patient data embeddings from the multiple heterogeneous data types taken from MIMIC-IV-ED and MIMIC-CXR. The embeddings are formed from numerical, categorical, textual, temporal, and image data.
4 Methods
The patient data from MIMIC-CXR and MIMIC-IV-ED for an exam are transformed into embeddings, which are used to prompt a multimodal language model to generate the findings and impression sections of the radiology report, as illustrated in Figure 2. Additionally, ‘Source’ embeddings differentiate the source of the data (e.g., the ‘chief complaint’ column from the triage table, the indication section, etc.), and time delta embeddings represent the time difference between an event and the exam. Standard embeddings, such as position and token embeddings, are also included. The patient data embeddings originate from three main groups: the tables of MIMIC-IV-ED; the report, images, and metadata of the current exam from MIMIC-CXR; and the patient’s prior exams (also originating from MIMIC-CXR). The prior exam and image embeddings are described in Section A and Subsection C.2, respectively.
4.1 Time, Position, & Source Embeddings
The ED information from MIMIC-IV-ED is typically recorded as discrete events, such as medications administered or vital signs measured, each with a specific timestamp. Events that occur closer to the time of the patient’s exam are generally more relevant for diagnostic purposes. To capture this, a time delta is calculated by subtracting the time of an event from the time of the exam. The exam time originates from MIMIC-CXR’s metadata table (Figure 3), whereas most of the MIMIC-IV-ED tables have event times for each row. As shown in Figure 3, the time delta is first converted to hours and then mapped using $1/\sqrt{\bm{\Delta}+1}$, assigning higher weights to events that occurred closer to the exam. The mapped time deltas are then passed through a feedforward neural network (FNN) defined as $f(\bm{\Delta}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$, where $\bm{W}_{\text{1}}\in\mathbb{R}^{1,2048}$, $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $f(\cdot)$ is the sigmoid linear unit (SiLU) activation function (Hendrycks and Gimpel, 2016), and $H$ is the hidden size of the decoder. This process generates the time delta embeddings, which are subsequently added to the embeddings of their respective sources. As shown in Figure 2, time delta embeddings are only applied to the prompt. Patient data from the current exam, such as the images, have a time delta of zero, while data from prior exams have a positive time delta.
The position embeddings are ordered by the time delta (Figure 3). This is due to the rotary position embeddings of the decoder; tokens that are closer together are given more importance. Hence, the smaller the time delta, the closer the embedding’s position is to the report token embeddings. Following Nicolson et al. (2024a), each unique patient data source is given its own source embedding. This includes the images, each report section, each table’s text column and value-category columns (described in the next section), and prior images and report sections.
4.2 Tabular Data
An example table and its conversion to embeddings is shown in Figure 3. To convert an exam’s tabular data to embeddings, columns were designated as value, category, text, or time columns. Value columns contained numeric data, while category columns contained categorical data. Datum from value and category columns were grouped by their time delta, with each group forming a feature vector. The feature vector initially consisted of zeros. Values and categories from the group were then used to set its values based on indices determined by a lookup table. For value columns, the lookup table determined the index where the numeric value was placed. For category columns, it determined which indices were activated (set to 1).
Next, the feature vector was passed through an FNN $f(\bm{X}_{i}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$ to form the embedding, where $\bm{X}_{i}\in\mathbb{R}^{|U_{C}|,|L_{i}|}$ are the grouped features, $\bm{W}_{\text{1}}\in\mathbb{R}^{|L_{i}|,2048}$ and $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $L_{i}$ is a lookup table, and $i$ designates the table. Each table has a unique FNN and lookup table. Rows for a value column always had a unique time, preventing multiple values from the same column in a group. We investigated alternatives to form the value-category embeddings in Section 5. The described framework was found to be the most efficient. Columns with a high cardinality were set as text columns. Text embeddings were formed via the decoder’s tokenizer and token embeddings. Text embeddings were given the time delta embedding from their respective row. The column designation for each table in Figure 1 is described in the Appendix B.
4.3 Report Section Embeddings
Here, we consider five sections of the radiology report: the findings, impression, indication, history, and comparison sections. The findings and impression sections serve as the ground truth to be generated. The remainder form part of the patient data embeddings. The indication section explains the reason for the exam, such as symptoms or suspected conditions. The history section provides relevant medical history, such as past conditions and treatments. The comparison section mentions any prior exams used to identify changes over time. These sections provide context that guides the interpretation of the exam, influencing the content of the findings and impression sections. The embeddings were formed via the decoder’s tokenizer and token embeddings. Of these, the history and comparison sections have not been investigated for CXR report generation. The comparison section was used only when prior exams were considered.
[…]
Images only
Images 26.00 29.24 58.87 24.10 12.24 14.35 24.34 6.33 272.4
Patient Emergency Department (ED) data (MIMIC-IV-ED)
Images + ED stays 26.10 29.47 60.65 24.17 12.39 14.52 24.50 6.36 273.4
Images + triage 26.46 31.27 63.06 24.29 12.32 14.66 24.58 6.44 278.9
Images + vital signs 26.47 31.72 63.39 24.32 13.16 14.61 24.74 6.47 274.7
Images + medicine reconciliation 26.86 31.37 63.98 24.52 12.77 14.90 24.85 6.60 343.5
Images + medicine administration 26.15 29.47 59.21 24.25 12.30 14.44 24.47 6.38 273.0
Patient additional radiology data (MIMIC-CXR)
Images + indication 26.94 32.13 65.43 24.74 14.16 15.19 25.16 7.02 279.5
Images + history 27.00 31.88 65.06 25.05 14.32 15.30 25.48 7.33 277.0
Images + metadata 26.34 29.63 59.55 24.37 12.40 14.55 24.50 6.43 273.4
Prior exams
Images + $h=1$ 26.98 31.42 63.98 24.65 12.65 15.11 25.03 6.78 558.9
Images + $h=1$ + comparison 26.76 31.55 64.20 24.42 13.36 15.03 24.82 6.74 563.4
Images + $h=2$ 26.67 30.48 61.27 24.53 13.60 14.94 24.85 6.72 810.6
Images + $h=2$ + comparison 26.20 30.19 61.24 24.05 12.43 14.80 24.55 6.58 815.0
Images + $h=3$ 26.47 29.96 59.95 24.14 12.90 14.94 24.66 6.65 1037.1
Images + $h=3$ + comparison 26.14 30.09 60.51 23.90 13.22 14.87 24.56 6.64 1041.5
All effective sources (no medicine administration, metadata, or comparison)
Images + effective sources ($h=0$) 27.11 32.23 64.80 25.07 14.48 15.15 25.40 7.07 365.0
Images + effective sources ($h=1$) 26.78 31.83 63.85 24.75 14.10 15.15 25.25 7.01 651.7
Ablation from Images + effective sources ($h=0$)
    - medicine reconciliation 26.78 32.81 65.60 24.84 14.44 15.21 25.33 7.19 293.9
    - ED stays 26.94 31.56 64.87 25.02 14.08 15.14 25.37 7.09 364.0
    - triage 27.15 32.45 65.18 25.15 14.80 15.27 25.54 7.25 358.5
    - vital signs 27.27 31.78 65.44 25.14 14.07 15.35 25.49 7.22 362.6
    - indication 26.89 31.25 64.65 24.99 13.87 15.07 25.39 7.00 357.9
    - history 26.96 31.87 64.02 24.86 14.60 15.10 25.24 7.04 360.3
    - time delta 27.17 32.11 65.10 25.18 14.64 15.24 25.54 7.16 365.0
4.4 Experiment Setup
Our multimodal language model, illustrated in Figure 2, is based on CXRMate-RRG24; it features a Llama decoder and the UniFormer as the image encoder. The training procedure for our model involved three stages: (1) initial training on the MIMIC-CXR training set using only images as input with Teacher Forcing (TF) (Williams and Zipser, 1989), (2) further training on the dataset described in Section 1 with the inputs detailed in Table 1, again using TF, and (3) reinforcement learning on the same dataset through self-critical sequence training (SCST) Rennie et al. (2017) (only for Table 2). Our evaluation metrics included three that capture the semantics of radiology reporting — RadGraph-F1 (RG), CheXbert-F1 (CX), and CXR-BERT (CB) — as well as five natural language generation metrics: BERTScore-F1 (BS), CIDEr (C), METEOR (M), ROUGE-L (R-L), and BLEU-4 (B4). Comprehensive details on the model architecture, training procedure, significance testing, and comparison methods are provided in Appendix C.
Table 2: Benchmark models on the test set described in Section 3 ($n=958$). Evaluation is on the findings section only. Underlined indicates statistical significance between the top two scores ($p<0.05$). In the ‘Train samples’ column, ‘images’ means the model generates reports per image, while ‘exams’ means a report generated per exam.
Model Train samples RG CX CB BS C M R-L B4
EMNLI Miura et al. (2021) 152 173 exams 32.8 28.9 66.6 24.4 19.4 17.1 28.1 8.9
CMN Chen et al. (2021) 270 790 images 25.3 24.3 49.4 19.7 16.9 15.1 26.4 7.6
TranSQ Kong et al. (2022) 368 960 images 29.8 30.4 62.3 20.4 14.9 17.6 22.6 7.9
RGRG Tanida et al. (2023) 166 512 images 23.2 22.8 37.9 23.4 7.6 12.4 21.1 5.4
CvT2DistilGPT2 Nicolson et al. (2023) 270 790 images 25.8 29.3 59.8 24.8 20.9 16.0 27.3 8.8
RaDialog Pellegrini et al. (2023) 276 778 images 26.8 38.4 60.7 26.2 14.6 14.7 25.4 6.9
MedXChat Yang et al. (2023) 270 790 images 22.6 13.1 21.3 19.3 9.8 14.3 23.2 7.0
CXR-LLaVA-v2 Lee et al. (2024) 193 513 images 20.7 20.7 44.1 23.6 5.2 11.3 19.9 2.7
CXRMate Nicolson et al. (2024a) 125 395 exams 28.8 33.9 71.3 30.5 22.4 17.7 28.1 9.7
CXRMate-RRG24 Nicolson et al. (2024b) 550 395 exams 30.4 31.2 58.2 31.0 20.6 16.7 27.5 9.1
Ours 76,398 exams 33.7 35.1 79.1 35.8 24.1 19.1 30.6 11.9
Figure 4: Case study demonstrating how incorporating a diverse set of patient data can aid with report generation.
5 Results & Discussion
The impact of different patient data sources on the performance of CXR report generation is summarised in Table 1. This analysis identifies which additional data sources enhance performance compared to using only images.
Significant improvements were observed by incorporating either the ED stays, triage, medicine reconciliation, or vital signs data from MIMIC-IV-ED dataset. Notably, the ED data markedly improved scores on the radiology report metrics (RG, CX, and CB). The medicine administration table did not significantly improve the scores overall, likely due to its infrequent occurrence in the exams (37%). (However, as shown in Table 4, it significantly improves performance when evaluated solely on exams that include a medicine administration table.) These findings demonstrate that ED patient data can enhance the diagnostic accuracy of CXR report generation.
Incorporating the indication or history section led to significant score improvements. This demonstrates the substantial influence these sections have on the findings and impression sections. Conversely, adding the metadata table did not result in significant score improvements, indicating it lacks valuable information for CXR report generation. While previous studies have established that the indication section boosts CXR report generation Nguyen et al. (2023), our findings demonstrate that the history section is equally important.
[…]
Next, we further trained ‘Images + effective sources (h = 0) - medicine reconciliation’ with reinforcement learning, as described in Subsection 4.4. This model, denoted as ‘Ours’ in Table 2, was compared to other benchmark CXR report generation models in the literature that included MIMIC-CXR in their training data. Despite having substantially fewer training samples than the other models, our model significantly outperformed them on CXR-BERT, BERTScore-F1, METEOR, ROUGE-L, and BLEU-4. This demonstrates the impact of incorporating a more comprehensive set of patient data on CXR report generation.
A case study is presented in Figure 4 demonstrating how a diverse set of patient data can impact report generation. Here, the first model is given the image only, and fails to identify key findings that the radiologist noted in their report. The second model is given the additional patient data available for this exam; the indication section and triage data. Hypoxia, as indicated by the low oxygen saturation (‘o2sat’), along with the elevated respiratory rate (‘resprate’) and systolic blood pressure (‘SBP’), are consistent with the physiological responses to pulmonary edema. Given this, the second model was able to identify the moderate pulmonary edema, echoing the radiologist’s findings.
Table 3 compares different methods for converting value and category columns into embeddings. This evaluation includes images, the triage table, and the medicine reconciliation table, as these tables contain multiple value and category columns. The aforementioned method of producing embeddings by grouping data from value and category columns (‘Grouped embeddings’) is compared to two other methods. The first is separate embeddings for each datum, where each value column datum is separately transformed using the previously described FNN, while each category column datum is converted to an embedding using a learnable weight matrix, akin to how token embeddings are produced (‘Separate embeddings’). The second method modifies ‘Separate embeddings’ by instead converting the value column data to text and using the decoder’s tokenizer and token embeddings (‘Values-to-text, categories-to-tokens’). The results indicate that the grouped embeddings method generally works best and is useful for encoding heterogeneous patient data for multimodal models.
6 Conclusion
This paper demonstrates the value of incorporating diverse patient data into automated CXR report generation. By integrating patient data from the MIMIC-CXR and MIMIC-IV-ED datasets, we have shown significant improvements in the diagnostic accuracy of generated radiology reports. Our empirical evaluation uncovers new sources of patient information that enhance CXR report generation, including data from ED stays, triaging information, aperiodic vital signs, medications, and the history section of radiology reports. We present specific methods to convert multimodal patient data into embeddings for a language model, encompassing numerical, categorical, textual, temporal, and image data. We encourage further research and experimentation using our released dataset splits, code, and model checkpoints to explore innovative methods for multimodal patient data integration, with the ultimate goal of enhancing diagnostic accuracy and patient care.
[…]
Embeddings CX RG CB BS
Images
Baseline 25.81 29.00 59.04 23.85
Images + triage + medicine reconciliation
Grouped embeddings 26.72 31.69 64.01 24.38
Separate embeddings 25.32 25.28 46.29 23.51
Values-to-text, categories-to-embeddings 26.46 30.70 58.62 24.58
7 Limitations
Despite the promising results demonstrated in this study, several limitations must be acknowledged. Firstly, the generalisability of our findings may be constrained by the datasets utilised, specifically MIMIC-CXR and MIMIC-IV-ED, which are derived from a single institution, the Beth Israel Deaconess Medical Center. This could introduce biases unique to the demographic and clinical practices of this institution, potentially limiting the applicability of our model to other healthcare settings with different patient populations or clinical workflows. Our reliance on these datasets is due to the fact that they are the only publicly available sources that link CXR exams with ED records.
Another limitation pertains to the completeness and quality of the patient data. Despite incorporating a wide range of data sources, the datasets still contain missing or incomplete information, which can affect model performance. For example, not all exams include a history section, and not all ED patient records have medicine administration details, leading to potential gaps in the data that the model can utilise. However, this reflects the nature of real patient records where issues of data quality and completeness are to be expected.
Our model’s architecture, while effective, has certain limitations. It struggles with large input sizes, especially when incorporating multiple prior exams, likely due to attention dilution. Future work should explore advanced attention mechanisms or hierarchical models to better manage large input sequences.
The interpretability of the model also poses a challenge. While our model shows improved diagnostic accuracy, the decision-making process within the multimodal language model remains a black box. Developing methods to enhance the interpretability and explainability of the model’s outputs would be beneficial, especially in clinical settings where understanding the rationale behind a diagnosis is critical.
Finally, while we provide a comprehensive set of metrics to evaluate our model’s performance, these metrics focus primarily on the diagnostic accuracy and quality of the generated reports. Broader evaluations considering clinical outcomes, such as the impact on patient management or reduction in radiologist workload, would offer a more holistic view of the benefits and limitations of CXR report generation models in general. Conducting such assessments could help to better understand the practical implications of deploying these models in a clinical setting.
In summary, while our study provides valuable insights into the integration of multimodal patient data for CXR report generation, addressing these limitations will be crucial for further advancements and broader adoption of such models in clinical practice. Future research should explore alternative architectures and training strategies, find alternative datasets to evaluate generalisability, improve model interpretability, and comprehensively assess the practical impact on patient care and radiologist workflow.
8 Ethical Considerations
In this research, we used real-world patient data from the MIMIC-CXR and MIMIC-IV-ED datasets. Since these datasets are de-identified, we consider privacy leakage risks to be minimal. Our method employs a language model to generate medical reports from patient data. However, we acknowledge that language models can exhibit bias and produce hallucinations, which may result in incorrect content in the generated reports.
Appendix A Prior exam embeddings
The images, findings section, and impression section from previous exams were considered. For prior exams, the time delta was positive, calculated by subtracting the time of the prior exam from the current exam. The images, findings section, and impression section from prior exams were given distinct source embeddings, separate from the current exam, to enhance differentiation. The comparison section from the current exam was also investigated, anticipating that references to prior exams in this section would prompt the decoder to reflect this in the generated report. We explored prior exams with a history size $h$ of up to three.
Appendix B Table column determination
The columns from the tables described in Figure 1 were given the following designations:
  • For the ED stay table, the patients ‘intime’ was used as the event time. Gender (e.g., ‘F’), race (e.g., ‘HISPANIC OR LATINO’), and arrival transport (e.g., ‘AMBULANCE’) were designated as category columns. The disposition column was not considered.
  • For the triage table, the ‘intime’ from the ED stay table was used. Temperature (e.g., ‘100.6’), heart rate (e.g., ‘93’), respiratory rate (e.g., ‘16’), O2 saturation (e.g., ‘94’), systolic blood pressure (SBP) (e.g., ‘110’), diastolic blood pressure (DBP) (e.g., ‘56’), and acuity (e.g., ‘2’) were designated as value columns. Pain (e.g., ‘6-9’ and ‘yes.’) and the chief complaint (e.g., ‘BILATERAL FOOT PAIN’) were designated as text columns.
  • The column designations for the vital sign table were identical to the triage table, except for the rhythm column (e.g., ‘Normal Sinus Rhythm’), which was treated as a category column. The vital signs table also had no chief complaint column and the ‘charttime’ column was used as the event time.
  • For the medicine reconciliation table, the ‘intime’ from the ED stay table was used as the event time, as it pertains to the patient’s medication history prior to the ED stay. The name column was designated as a text column, while the gsn, ndc, etc_rn, and etccode columns were designated as category columns. The etcdescription column was not considered, as it is a descriprion of the etccode column.
  • For the medicine administration (pyxis) table, ‘charttime’ was used as the event time. The med_rn, name, gsn_rn, and gsn columns were all treated as category columns. The name column for the medicine reconciliation column did not have as high of a cardinality as the name column from the medicine reconciliation column, allowing it to be considered as a category column.
  • For the metadata table, the ‘PerformedProcedureStepDescription’, ‘ViewPosition’, ‘ProcedureCodeSequence_CodeMeaning’, ‘ViewCodeSequence_CodeMeaning’, and ‘PatientOrientationCodeSequence_CodeMeaning’ columns were considered, and designated as category columns.
Appendix C Experiment setup
C.1 Metrics
CheXbert-F1 (Smit et al., 2020), RadGraph-F1 Delbrouck et al. (2022), BLEU-4 (Papineni et al., 2001), and BERTScore-F1 (roberta-large_L17_no-idf_rescaled) (Zhang et al., 2020) have been found to correlate with radiologists’ assessment of reporting (Yu et al., 2023) and were a part of our evaluation. Additionally, we include CXR-BERT Boecking et al. (2022); Nicolson et al. (2024a), CIDEr (Vedantam et al., 2015), METEOR Banerjee and Lavie (2005), and ROUGE-L (Lin and Hovy, 2003) as part of our evaluation. CheXbert-F1, RadGraph-F1, and CXR-BERT were intended to capture the clinical semantic similarity between the generated and radiologist reports, while BERTscore-F1 was intended to capture general semantic similarity. Finally, CIDEr, METEOR ROUGE-L, and BLEU-4 were intended to capture the syntactic similarity between the generated and radiologist reports.
For the models in Table 2 that generate a report for each image in an exam, the average score was taken across all reports for an exam. Following this, the final average score was computed across all exams for both models that generate a report per image and those that generate a report per exam.
For CheXbert, the macro-averaged F1 was computed between the 14 CheXbert observations extracted from the generated and radiologist reports. “No mention”, “negative”, and “uncertain” were considered negative, while “positive” was considered positive. Here, the true positives, false positives, and false negatives were averaged over the reports of each exam for the models that generate a report per image.
We also perform statistical testing; first, a Levene’s test was conducted to reveal if the variances across model scores was homogeneous or not. If the assumption of equal variances was upheld, a one-way ANOVA was conducted to determine if there was a significant difference between models. Finally, pairwise Tukey-HSD post-hoc tests were used for pairwise testing. If the assumption of equal variances was violated, a one-way Welch’s ANOVA was conducted to determine if there was a significant difference between models. Finally, Games-Howell post hoc tests were used for pairwise testing. A $p$-value of 0.05 was used for all significance testing. Statistical testing was not performed for CheXbert, as it is a classification metric.
C.2 Model
Our model is illustrated in Figure 2; following Nicolson et al. (2024b), we utilised UniFormer as the image encoder (in particular, the $384\times 384$ base model warm started with its token labelling fine-tuned checkpoint) Li et al. (2023). The image embeddings are formed by processing each image in the exam separately with the image encoder and then projecting its last hidden state to match the decoder’s hidden size using a learnable weight matrix. Each image was resized using bicubic interpolation so that its smallest side had a length of 384 and its largest side maintained the aspect ratio. Next, the resized image was cropped to a size of $\mathbb{R}^{3\times 384\times 384}$. The crop location was random during training and centred during testing. Following Elgendi et al. (2021), the image was rotated around its centre during training, where the angle of rotation was sampled from $\mathcal{U}{[{-5^{\circ},5^{\circ}}]}$. Finally, the image was standardised using the statistics provided with the UniFormer checkpoint. A maximum of five images per exam were used during training. If more were available, five were randomly sampled uniformly without replacement from the exam.
Again following Nicolson et al. (2024b), we employed the Llama architecture for the decoder, which is notable for features such as its rotary positional encoding (RoPE), root mean square normalisation (RMSNorm), and SwiGLU activation function Touvron et al. (2023). A byte-level byte pair encoding tokenizer (Wang et al., 2020) was trained with a vocabulary size of 30 0003000030\,00030 000. It was trained on the findings, impression, indication, and history sections (not the comparison section) of the entire MIMIC-CXR training set, as well as the ‘pain’ and ‘chiefcomplaint’ columns from the triage table, the ‘name’ column of the medicine reconciliation table, and the ‘pain’ column from the vital signs table (from the entire MIMIC-IV-ED dataset). Newline, tab, repeated whitespaces, and leading and trailing whitespaces were removed from any text before tokenization.
The hyperparameters of the Llama decoder were six hidden layers, a hidden size of 768, 12 attention heads per layer, and an intermediate size of 3 07230723\,0723 072. The maximum number of position embeddings was set to 2048 to accommodate all the patient data embeddings and the report tokens. The maximum number of tokens that could be generated was set to 256, which was also the limit for the radiologist reports during training. During testing, a beam size of four was utilised. The Llama decoder allows a custom attention mask to be provided in current implementations. [^4] https://huggingface.co/blog/poedator/4d-masks This enabled non-causal masking to be utilised for the prompt and causal masking for the report token embeddings, as shown in Figure 5. This ensured that the self-attention heads were able to attend to all of the patient data embeddings at each position.
Figure 5: Attention mask for the decoder. Non-causal masking was used for the patient data embeddings and causal masking for the report token embeddings.
[…]
C.3 Training
Three stages of training were performed. Each stage used AdamW (Loshchilov and Hutter, 2022) for mini-batch gradient descent optimisation, where training and evaluation was performed on a 94GB NVIDIA H100 GPU. The three stages were as follows:
  1. 1.
    Teacher forcing (TF) Williams and Zipser (1989) was performed on the MIMIC-CXR dataset with only the images for an exam as input, and exams that contained both a findings and impression section. This gave a training/validation split of 232 853232853232\,853232 853/1 83718371\,8371 837 images, 125 416125416125\,416125 416/991 exams, and 57 1015710157\,10157 101/436 patients. Training was performed with an initial learning rate of 5e-5, a mini-batch size of 8, a maximum of 32 epochs, and with float16 automatic mixed precision. All model parameters were trainable during this stage. The validation macro-averaged CheXbert-F1 was the monitored metric for checkpoint selection. This stage was necessary, as the language model struggled to generate reports from multiple sources without prior learning.
  2. 2.
    TF on the dataset described in Section 3 with the inputs described in Table 1. The training strategy was identical to the previous stage, except that a maximum of 16 epochs was performed, and the image encoder’s parameters were frozen (except for its projection). The models featured in Table 1 were trained using only the first two stages.
  3. 3.
    Reinforcement learning using self-critical sequence training (SCST) Rennie et al. (2017) with CXR-BERT and BERTScore as the reward (each weighted with 0.5) was performed in the final stage of training. The sample report for SCST was generated with top-k sampling ($k=50$). Training was performed with an initial learning rate of 5e-6, a mini-batch size of 32, a maximum of 24 epochs, and with float32 precision. The image encoder’s parameters were frozen during this stage (except for its projection). The validation BERTScore-F1 was the monitored metric for checkpoint selection, as it helped to select checkpoints less prone to repetitions. This stage of training was only applied to the best model from Table 1, ‘Images + effective sources (h = 0) - medicine reconciliation’, with the results presented in Table 2. This model had 161 185 728161185728161\,185\,728161 185 728 parameters.
C.4 Comparison Models
The generated reports for the models in Table 2 were attained as follows:
CXRMate-RRG24 was trained on five datasets, including MIMIC-CXR. RGRG was trained on the ImaGenome dataset derived from MIMIC-CXR — which may have some overlap with our test set.
[…]
Inputs RG CX CB BS
Images 26.24 28.36 57.17 24.33
Images + medicine administration 26.95 28.53 58.94 24.93
Appendix D Ancillary results
In Figure 6, the F1-scores for each CheXbert label are shown. The ‘Images + effective sources ($h=0$) - medicine reconciliation’ model from Table 1 improves performance across all labels compared to the ‘Images’ model. This suggests that incorporating ancillary data from MIMIC-IV-ED and MIMIC-CXR provides a general improvement, rather than benefiting any specific pathology.
Further improvements are seen when training the ‘Images + effective sources ($h=0$) - medicine reconciliation’ model with SCST (i.e., our model from Table 2) for most pathologies. However, there are performance decreases for ‘enlarged cardiomediastinum’, ‘lung lesion’, ‘pneumothorax’, and ‘fracture’. This might be due to these pathologies being underrepresented in the MIMIC-CXR dataset, leading the model to optimise for more common pathologies during SCST.
The results for exams that include a medicine administration table are show in Table 4. Adding the medicine administration table produced a significant improvement in the scores, indicating that it should be considered if available.